OpenStudy (anonymous):

please answer this: How will you determine the equation of the line if the given are two points of the line? i really need the answer... thank you

OpenStudy (anonymous):

first find the slope, m=(y2-y1)/(x2-x1)

OpenStudy (anonymous):

then write it in slope intercept form y-y1=m(x-x1)

Directrix (directrix):

What if I make up two points and work the problem I made up? Will that help?

OpenStudy (anonymous):

okey....

Directrix (directrix):

Points ( 2, -1 ) and ( -4, -5 ) lie on a line. Find the equation of the line that contains then. Leave your answer in y = mx + b form.

Directrix (directrix):

First, get the slope. Think of slope as change in y over change in x. m designates the slope of a line and has formula [y2 - y1] / x2 - x 1 ] . The y2 is not y^2 but what we call y sub 2 because the 2 is a subscript which is hard to type.

Directrix (directrix):

One of our points will be (x1, y1) and the other (x2, y2) It does not matter which is which.

Directrix (directrix):

Let (x1,y1) be (2, -1) and let (x2, y2) be (-4, -5) Questions so far?

OpenStudy (anonymous):

what is y=mx+b

Directrix (directrix):

That is the form we will leave the equation of our line in. So, if we get something like 2x + 7y = 2/3, then that's not y = mx+b form.

OpenStudy (anonymous):

ah okey

Directrix (directrix):

You don't have to leave the equation in that form unless instructed to do so. Most books do request that or standard form.

Directrix (directrix):

Focus on cranking out this slope right now.

OpenStudy (anonymous):

okey

Directrix (directrix):